OpenStudy (anonymous):

Ms. Millie showed the graphs of the functions y=log2x and y=log4x to her students. Which conclusion is incorrect? A. The x-intercept of each graph is 0 B. The graphs never intersect the y-axis C. The domain of each function is {x:x>0} D. The range of each function is {y: all real numbers}

OpenStudy (anonymous):

I am not quite sure, but I think it is the letter B?

OpenStudy (jdoe0001):

graph them, look at them

OpenStudy (anonymous):

okay, by definition, what is log(a)?

OpenStudy (anonymous):

I don't know how to put the 4 in TL-84 calculator

OpenStudy (anonymous):

Log is a power

OpenStudy (anonymous):

log(a)=x is the solution to the equation e^(x)=a.

OpenStudy (anonymous):

is e you mean lnx?

OpenStudy (anonymous):

e is exponential.

OpenStudy (anonymous):

Oh, okay...

OpenStudy (anonymous):

So if you substitute in your question x=0, you get log0. What is log 0?

OpenStudy (anonymous):

e^o=1

OpenStudy (anonymous):

um, i have confused you a little bit, the two x's are different.

OpenStudy (anonymous):

in your problem you're looking for log x=a log 0=a e^(which power)=0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, so it should be clear that there is no power of e that will give you 0. So the answer to your question is A.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

I'm more concerned about you understanding it, than getting the answer. Is everything clear?